OpenStudy (anonymous):

1.Simplify 4 x all over 9 plus quantity x minus 10 over 9 (1 point)

OpenStudy (anonymous):

4x/(9+((x-10)/9)) is that what you are saying?

OpenStudy (anonymous):

Look at the rationals in your example. They are divided by 9. So multiply the non-rationals by 9/9 (which is 1). That would be x. Now the example becomes 4x/9 + 9x/9 -10/9 They all have the same denominator, so we can group them all over 9. (4x + 9x - 10)/9 combine like terms. (13x - 10)/9 For the other problem, the first thing to do is distribute the negative sign across the parentheses. (x^2 + 2x/12x + 54) - 3 + x/8x - 36 = (x^2 + 2x/12x + 54) - 39 + x/8x Now the parentheses on the first part don't matter any more and you can combine like terms. x^2 + 2x/12x + x/8x + 54 - 39 = x^2 + 2x/12x + x/8x + 15 On the 2 rational expressions, note that both of them have x over x. Anything divided by itself becomes 1. In a way, it sort of disappears. 2x/12x = 2/12 * (1) = 2/12 = 1/6 Then x/8x becomes 1/8. You have to cross-multiply to add 1/8 and 1/6 which will become 14/48 = 7/24. Now your equation reduces to x^2 + 15 and 7/24.

OpenStudy (anonymous):

Thank you so much! Could you answer a couple more for me? This is the only section of my lesson I don't understand

OpenStudy (anonymous):

2.Simplify quantity t squared plus 6 t over the quantity t minus 1 end quantity minus 7 over quantity t minus 1